The area of a square with side 6 cm is one-sixth of the area of a rectangle whose length is six times its breadth. What is the perimeter of the rectangle?
Answer by mananth(16946)

Side of square = 6 cm
area = 6*6 = 36 cm^2
..
this is 1/6 of rectangle area
rectangle area = 6 times area of square
36*6 = 216 cm^2
..
Area of rectangle = L*W
let width be x cm
length = 6x
6x*x = 216
6x^2= 216
/6
x^2= 216/6
x^2=36
take the square root
x = 6 cm
..
m.ananth@hotmail.ca